OpenStudy (anonymous):

car was going 70mph then when it stop it went 177ft. how much will car go if it stops at 15 mph and 130 mph

OpenStudy (anonymous):

it was wrong.. the stopping distance is proportional to the square of the velocyity

OpenStudy (anonymous):

oh!!!!

OpenStudy (anonymous):

you didnt say that so we have s = k * v^2 where k is a constant we find k by plugging in the values v=70 and s = 177 177 = k * 70^2 k = 177/70^2 = 0.0361 so the formula relating stopping distance and velocity is s = 0.0361v^2 so plug in v=15 s = 0.0361 * 15^2 = 8 ft to find s for 130 mph plug in v=130

OpenStudy (anonymous):

does feet and mile matter?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

if u convert to ft you'll get the same answer but the value of k will differ - and you'll be dealing with BIG numbers

OpenStudy (anonymous):

thank you master